OpenStudy (anonymous):

Estimate the area of a square floor if each side is 12 ft 9 in long. Estimate the area of a square floor if each side is 12 ft 9 in long. @Mathematics

OpenStudy (anonymous):

23409 in^2

OpenStudy (anonymous):

169

OpenStudy (anonymous):

because 13 *13 =169

OpenStudy (radar):

This problem is to test your understanding of units. Note that the dimension is given in two units, feet and inches. You know that area is the product of length times width, of course in a square length is the same as width. But how do you multiply when two different units. There is two ways to attack this problem. you could convert the dimension totally to inches as the first poster did, your answer will then be in square inches. Or you can convert it to feet, using feet your answer will be in square feet, which is more common when figuring flooring, carpeting etc. Step 1. Convert the 9 inches to feet 9/12=3/4=.75 ft. The dimension is therefore 12.75. Area = 12.75 X 12.75 = 162.5625 sq ft.

OpenStudy (anonymous):

but you need to estimate it.

OpenStudy (anonymous):

ooh , i guess you did. but really closely.

OpenStudy (radar):

Dividing abdul shabeer's answer by 144 would convert the square inch answer to square ft. 23409/144=162.5625

OpenStudy (radar):

Yes I see it merely calls for an estimate. You can estimate as close as you want to!

OpenStudy (radar):

To estimate I would use 12 x 13 = 156 That will get you closer than 13 X13 = 169

OpenStudy (anonymous):

ok